• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M. Stevens

    Log entry time 15:11:17 on March 27, 2007

    Entry number 197551

    keyword=Spectrometer magnets functionally test o.k.

    Performed polarity reversal, lead flow interlock and helium level interlock tests on all magnets. Ramped left spectrometer to 4 GeV with no problems. Ramped right spectrometer to about 1.5 GeV before I had to ramp it down to facilitate other work. The right spectrometer magnets also worked fine, except for a minor casualty- the right Q1 local control panel lcd display has died, but the control panel is otherwise functional.

    A copy of this log entry has been emailed to: folts@jlab.org